a large bucket holds 3L more than twice as much as a small bucket. it took 2 small buckets and 5 large buckets to fill a 63L. how much does a large one hold

Answer by checkley79(3341)

L=3+2S
2S+5L=63
2S+5(3+2S)=63
2S+15+10S=63
12S=63-15
12S=48
S=48/12
S=4 SMALL BUCKETS SIZE.
L=3+2*4
L=3+8
L=11 LARGER BUCKET SIZE.
PROOF:
2*4+5*11=63
8+55=63
63=63